# Volumetric Surfaces: Representing Fuzzy Geometries with Multiple Meshes

### v1

@@ 5.1 Ablations
We ablate all crucial aspects of our method and show results on both implicit (Table 3) and explicit (Table 4) geometry phases. We run our full model:
  1. 1.
    without view-dependent surface transparency; we observe increased model expressivity as shown by the improved image quality metrics.
  2. 2.
    without curvature loss during $k$-SDF training ($\lambda_{s}=0.0$), surfaces can reconstruct high-frequency details, but the image quality of the baked representation worsens. This occurs because the simplified post-baking mesh poorly aligns with the implicit one. Conversely, enabling curvature loss results in $k$-SDF reconstructing smoother surfaces with meshes that closely align with their implicit counterparts, maintaining a low triangle cost. Additionally, this demonstrates how our high-capacity appearance modeling compensates for missing geometric details.
  3. 3.
    without transparency attenuation (Section 4.2) at grazing angles (no $\alpha_{w}$). We observe how it significantly reduces rendering errors, particularly at object boundaries.
  4. 4.
    with constant support surface offsets ($\Delta o$), which are not trained and remain fixed at their initial values. We observe how results worsen significantly. Training offsets allows spacing to adapt optimally to each scene.
Visualizations of the ablations are provided in our supplementary material.
Table 3: Ablation results on $5$-SDF averaged over Shelly dataset scenes [47]. See Section 5.1 for explanations.
Ablation PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
5-SDF, full 32.046 0.964 0.130
1) w/o view-dep. $\alpha$ 31.753 0.962 0.131
2) w/o curvature $\mathcal{L}_{\text{s}}$ 33.016 0.971 0.117
3) w/o $\alpha_{\text{w}}$ 32.106 0.966 0.130
Table 4: Ablation results on $5$-Mesh (baked, real-time rendering assets) averaged over Shelly dataset scenes [47]. See Section 5.1.
Ablation PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
5-Mesh, full 34.098 0.979 0.114
1) w/o view-dep. $\alpha$ 32.713 0.975 0.120
2) w/o curvature $\mathcal{L}_{\text{s}}$ 33.412 0.980 0.110
3) w/o $\alpha_{\text{w}}$ 33.963 0.980 0.113
4) w. const. $\Delta o$ 30.092 0.950 0.129
Table 5: Results on the khady scene from Shelly [47]. Sweep over the number of meshes and target neural textures resolution. Mixed resolutions are described in Section 4.5. Increasing neural texture resolution improves reconstruction quality up to $1024^{2}$, but performance declines beyond that. Modeling all textures at maximum resolution ($2048^{2}$) is counterproductive, as results show that a mixed-resolution approach is better in terms of image quality and memory footprint.
PSNR $\uparrow$
Method mixed 20482 10242 5122 2562
5-Mesh 29.82 29.82 29.85 29.76 29.48
7-Mesh 29.91 29.91 29.95 29.87 29.58
9-Mesh 29.93 29.90 29.94 28.87 29.59
@@ 5.2 Limitations
Experiments on the NeRF-Synthetic [25] dataset, which lacks fuzzy geometries, show only marginal improvements over PermutoSDF. This is likely due to increased view-dependency capacity, which tends to overfit training views and reduce test view quality. Similarly, our model excels in densely observed scenes but underperforms when training views are sparse, as our representation sacrifices multi-view consistent geometric complexity for view-dependent effects that do not generalize well to test views.” A solution explored in [44] involves initially training a robust model (e.g., NeRF-like) and then distilling its reconstruction by using renders from randomly generated cameras as the training set for our method. The handling of thin structures remains challenging due to the underlying SDF geometry representation.
@@ 6 Conclusion
We presented Volumetric Surfaces, a multi-layer mesh representation for real-time view synthesis of fuzzy objects on a wide range of devices. Our method renders faster than state-of-the-art volume-based approaches, while being significantly more capable at reproducing fuzzy objects than single-surface methods. In future work, we would like to develop a single-stage, end-to-end training procedure that directly produces real-time renderable assets.
Acknowledgments: Plushy blender object by kaizentutorials. S.Esposito acknowledges travel support from the European Union’s Horizon 2020 research and innovation program under ELISE Grant Agreement No 951847.
@@ 7 Supplementary
In this supplementary material, we present additional visualizations (Section 7.1 and Section 7.2), comprehensive per-scene results (Section 7.3), and a deeper analysis of performance on fully solid geometries (Section 7.4).
7.1 Transparency Attenuation
We introduced transparency attenuation in Section 4.2 to reduce visual artifacts at object boundaries. Figure 8, cropped from our ablation experiments, highlights its significance in our method.
Refer to caption Refer to caption
$a$) without $\alpha_{\text{w}}$ $b$) with $\alpha_{\text{w}}$
Figure 8: Rendering error crop without (a) and with (b) transparency-decay, resulting in a 2.13 dB PSNR gain. Scene from the Shelly dataset [47].
7.2 Additional Visualizations
We provide additional qualitative comparisons on our evaluation scenes of the DTU [16] dataset. Additionally, we provide visualizations of per-surface rendering before alpha blending (Figure 10 and Figure 11) to illustrate how each layer, based on its position and opacity, contributes with its view-dependent appearance model to the final image.
7.3 Per-scene Results
Table 6, Table 8 and Table 8 present the per-scene values that are averaged in Table 1.
7.4 Fully Solid Scenes
Although not our targeted use case, we tested our method on the fully solid scenes of the NeRF-Synthetic dataset [25] (which does not contain any fuzzy objects). In these scenes, our method outperforms PermutoSDF [35] but lags behind other baselines, as mentioned in Section 5.2. We approach fuzzy surface modeling not by emphasizing intricate geometric details but by increasing the number of samples with optimized spatial distribution, as these factors are crucial for the accurate representation of such effects. Therefore, our method promotes smoother surfaces, leading to smoothed angles and overly simplified geometry in under-observed areas (Figure 13). Fully solid scenes could be optimally modeled as a single surface. However, SDF-based methods struggle in handling thin structures, with the optimization failing to extract them reliability (BakedSDF [51], BOG[34]). Our surfaces smoothing, combined with view-dependent transparency, results in thin structures being reconstructed as view-dependent effects (Figure 12). Consequently, our model tends to overfit training views, yielding a larger gap between training and test views quality (Table 9).
Refer to caption Refer to caption
PermutoSDF [35] 7-Mesh (ours)
Refer to caption Refer to caption
3DGS [17] Ground truth
Refer to caption Refer to caption
PermutoSDF [35] 7-Mesh (ours)
Refer to caption Refer to caption
3DGS [17] Ground truth
Figure 9: Scenes from the DTU dataset [16].
@@ 7.4 Fully Solid Scenes
d) Surface colors (RGB).
Figure 10: 7-Meshes render buffers. Layers order: left to right is inner to outer. Individual layer colors and alpha buffers are blended as described in Section 4.2.
@@
d) Surface colors (RGB).
Figure 11: 7-Meshes render buffers. Layers order: left to right is inner to outer. Individual layer color and alpha buffers are blended as described in Section 4.2.
Table 6: Per-scene results for baselines.
PermutoSDF [35] 3DGS [17]
Shelly [47] fernvase 28.424 0.953 0.078 34.821 0.986 0.040
horse 34.677 0.993 0.040 41.451 0.997 0.038
khady 26.219 0.879 0.226 30.545 0.924 0.187
kitten 30.913 0.971 0.093 38.169 0.991 0.050
pug 29.480 0.953 0.168 35.959 0.983 0.089
woolly 29.387 0.949 0.167 31.710 0.969 0.130
Custom (ours + [38]) hairy_monkey 33.672 0.977 0.194 37.672 0.990 0.142
plushy 32.940 0.945 0.192 37.018 0.975 0.153
DTU [16] (scans 83, 105) dtu_scan105 34.783 0.985 0.124 35.505 0.984 0.102
dtu_scan83 37.842 0.991 0.072 40.615 0.994 0.070
Table 7: Our per-scene results.
@@
Shelly [47] fernvase 32.261 0.984 0.068 33.518 0.987 0.065
horse 38.175 0.997 0.039 39.668 0.998 0.035
khady 29.645 0.937 0.198 29.825 0.939 0.197
kitten 35.681 0.990 0.080 36.739 0.992 0.077
pug 33.569 0.982 0.141 34.014 0.985 0.138
woolly 30.118 0.972 0.179 30.825 0.976 0.172
Custom (ours + [38]) hairy_monkey 35.433 0.983 0.182 35.894 0.985 0.180
plushy 34.252 0.956 0.168 34.944 0.964 0.163
DTU [16] (scans 83, 105) dtu_scan105 34.607 0.979 0.123 34.875 0.980 0.115
dtu_scan83 37.878 0.989 0.066 38.132 0.990 0.063
Table 8: Our per-scene results.
@@
Shelly [47] fernvase 34.420 0.990 0.063 34.531 0.991 0.063
horse 39.896 0.998 0.034 39.120 0.998 0.035
khady 29.918 0.940 0.197 29.927 0.942 0.198
kitten 37.022 0.993 0.076 36.946 0.993 0.076
pug 33.984 0.984 0.139 33.994 0.984 0.140
woolly 30.746 0.977 0.167 30.799 0.976 0.167
Custom (ours + [38]) hairy_monkey 36.081 0.986 0.177 36.130 0.987 0.175
plushy 35.129 0.967 0.162 35.301 0.969 0.160
DTU [16] (scans 83, 105) dtu_scan105 35.331 0.983 0.112 35.191 0.982 0.111
dtu_scan83 37.918 0.991 0.062 38.461 0.990 0.062
Table 9: Results averaged across test scenes. In solid scenes, our method outperforms PermutoSDF (see Figure 12) but lags behind other baselines. We explain this behavior in Section 7.4. Methods marked with a $\star$ show results taken from original papers, as code is unavailable. MobileNeRF and Instant-NGP results on Shelly from AdaptiveShells. PermutoSDF trained until densities are fully peaked ($\phi_{\beta_{3}})$. Metrics not provided by a baseline are denoted with “—”.
@@
3DGS [17] 36.763 0.991 0.030 33.227 0.981 0.037
Instant-NGP [26] 33.180
PermutoSDF [35] 29.311 0.975 0.057 28.052 0.966 0.065
AdaptiveShells 31.840 0.957 0.056
QuadFields 31.000 0.952 0.069
MobileNeRF [5] 30.900 0.947 0.060
3-Mesh 32.255 0.982 0.061 28.366 0.957 0.085
5-Mesh 33.077 0.985 0.056 28.644 0.958 0.083
7-Mesh 33.159 0.985 0.057 28.748 0.959 0.083
9-Mesh 33.044 0.985 0.058 28.656 0.959 0.084
Refer to caption Refer to caption
PermutoSDF [35] 7-Mesh (ours)
Refer to caption Refer to caption
PermutoSDF [35] 7-Mesh (ours)
Figure 12: Top row: train view; bottom row: test view. Scene from NeRF-Synthetic [25].
Figure 13: Surface normals. Scene from NeRF-Synthetic [25]

### v2

@@ 5.1 Ablations
We ablate all crucial aspects of our method and show results on both implicit geometry and fully baked phases in Table 4. We run our full model: 1) Without view-dependent surface transparency, we observe increased model expressivity as shown by the improved image quality metrics. 2) Without curvature loss during $k$-SDF training ($\lambda_{s}=0$), surfaces can reconstruct high-frequency details, but the image quality of the baked representation worsens. This happens because the post-baking mesh poorly aligns with the implicit one. Enabling curvature loss pushes $k$-SDF to reconstruct smoother surfaces with meshes that better match their implicit counterparts while keeping the triangle cost low. Moreover, this highlights how our high-capacity appearance model compensates for missing geometric details. 3) Without transparency attenuation (Section 4.2) at grazing angles (no $\alpha_{w}$), rendering errors increase, especially at object boundaries. 4) With fixed (not trained) support surfaces offsets ($\Delta o$). We observe a significant decline in results, as training offsets is essential for adapting spacing optimally to each scene. 5) Initializing support surfaces outside the main SDF. We observe how this leads geometry to extend beyond the object silhouette. While training views can compensate through learned view-dependent transparency, test views suffer from degraded generalization. In contrast, initializing surfaces inside biases them to be tighter, preventing unwanted expansion. This happens because the reconstructed main surface is typically conservative, forming an outer shell that encloses the scene content, including fuzzy geometry (Figure 1).
@@ 5.2 Limitations
Textured shells [27, 26, 23] exhibit artifacts at grazing angles, especially when test views fall outside training coverage or model capacity is limited. Increasing shell count mitigates this but raises memory and computation costs, particularly during reconstruction. Artist-designed extruded textured fins [26] can address these artifacts, though learning this component remains challenging and is left for future work. Our model performs well in densely observed scenes but struggles in sparsely sampled ones. When under-constrained, it tends to explain observations through view-dependency rather than multi-view consistent geometry, leading to poorer generalization in test views. A solution explored by Wan et al. [53] trains a robust model (e.g., NeRF-like) and distills its reconstruction using renderings from randomly generated cameras as the training set. Handling thin structures remains challenging due to the limitations of the underlying SDF geometry representation. Advantages on fully solid surfaces are also marginal; see the supplement for further details.
@@ 6 Conclusion
We presented Volumetric Surfaces, a multi-layer mesh representation for real-time view synthesis of fuzzy objects on a low-power laptops and smartphones. Our method renders faster than state-of-the-art volume-based approaches, while being significantly more capable at reproducing fuzzy objects than single-surface methods. For future work, we aim to develop a single-stage, end-to-end training procedure that directly generates real-time renderable assets.
Acknowledgments
Plushy Blender object by kaizentutorials. Stefano Esposito acknowledges travel support from the European Union’s Horizon 2020 research and innovation program under ELISE Grant Agreement No. 951847.
@@ Acknowledgments
In this supplementary material, we provide additional architectural and technical details (Section S1), further visualizations (Section S2), comprehensive per-scene results (Section S3), and an in-depth analysis of performance on fully solid geometries (Section S4).
S1 Additional Technical Information
$\beta$ scheduling details: During training, the $\beta$ parameter is controlled by the scheduling of $v$ as $\beta=e^{10v}$. During the main surface training phase, $v$ linearly transitions from $v_{1}=0.3$ to $v_{2}=0.7$. During the training of $k$-SDF, it further progresses from $v_{2}=0.7$ to $v_{3}=1.0$. At $\beta_{2}$, the logistic distribution standard deviation is approximately $0.001$. We use this value to initialize offsets as constants ($\Delta o$). By the end of implicit surface training ($\beta_{3}$), it decreases to $0.00008$, resulting in fully peaked densities.
$k$-SDF Architecture: We encode 3D points using the trainable positional encoding from Rosu and Behnke [44], followed by a small MLP with three layers of 32 features each. Hidden layers employ GELU activations, while the final layer uses a linear activation to output the signed distance $d$ (our main SDF) and a geometric feature vector $\mathbf{z}$. We predict relative offsets using tiny MLP heads (a single layer with 32 units) with independent parameters, taking only $\mathbf{z}$ as input. This ensures that model complexity scales with the number of surfaces. To enforce the sign of the predicted offset, we apply a softplus activation multiplied by the desired sign. Finally, we compute the final ordered offsets by performing a cumulative sum over the predicted relative offsets, separately for negative and positive values.
Volumetric Appearance Architectures: We model RGB and transparency as two networks with identical architectures, differing only in their output dimensions (3 for RGB and 1 for transparency). Both models encode 3D points using the trainable positional encoding from Rosu and Behnke [44], followed by an MLP with three layers of 128, 128, and 64 features, respectively. Its input consists of the encoded position, a spherical harmonics encoding (with degree 3) of the view direction $\mathbf{v}$, the normal vector of the rendered SDF $\mathbf{n}$ and the geometric feature vector $\mathbf{z}$ predicted by $k$-SDF. Normals are computed as the normalized gradients of the SDFs; gradients are computed with finite differences, $\epsilon=10^{-4}$. Hidden layers use GELU activations, while the final layer applies a Sigmoid activation to produce outputs in the range $[0,1]$.
Neural Textures Architecture: During the mesh texturing phase, we use separate neural texture models for RGB and transparency per mesh. We encode 2D UV coordinates using the trainable positional encoding from Müller et al. [35], followed by a small MLP with two layers of 64 features each. Hidden layers use ReLU activations, while the final layer applies a linear activation to output per-channel spherical harmonics (SH) coefficients of degree 3, which are then decoded with view direction $\mathbf{v}$.
Points sampling: During volumetric rendering the number of uniformly sampled points per ray in the foreground area of the scene is 64. On top of these, 32 points are added with importance sampling. Additionally, if a scene is unbounded (e.g. DTU), we sample 32 additional points in contracted space [3]. Rays batch size is defined w.r.t. a target number of sample points which is up to a maximum of $512\times 64\times 32$ points.
Data handling: We use MVDatasets [11] to load datasets, manage training loop pixel iterators, and perform ray casting.
@@
d) Blending weights (contributions).
e) Surface colors (RGB).
Figure S1: Visualization of render buffers from our 5-Mesh model. Layers order: left to right is inner to outer. Individual layer color and alpha buffers are blended as described in Section 4.2. Results on the khady scene from Shelly [56].
Refer to caption Refer to caption
$a$) without $\alpha_{\text{w}}$ $b$) with $\alpha_{\text{w}}$
Figure S2: ($a$) Rendering error crop (averaged over color channels) without and ($b$) with transparency-decay, resulting in a 2.13 dB PSNR gain. Scene from the Shelly [56].
@@
d) Blending weights (contributions).
e) Surface colors (RGB).
Figure S3: Visualization of render buffers from our 7-Mesh model. Layers order: left to right is inner to outer. Individual layer colors and alpha buffers are blended as described in Section 4.2. Results our custom plushy scene.
S2 Additional Visualizations
We provide additional qualitative comparisons on our evaluation scenes of the DTU [21] dataset. Additionally, we provide visualizations of per-surface rendering before alpha blending (Figure S1 and Figure S3) to illustrate how each layer, based on its position and opacity, contributes with its view-dependent appearance model to the final image. Finally, we visualize results from Table 2. Figure S5 presents a qualitative comparison between a render of our 7-Mesh model, 3DGS [24], and 3DGS-75K. Figure S7 compares our 5-Mesh model to MobileNeRF [7].
S2.1 Transparency Attenuation
We introduced transparency attenuation in Section 4.2 to reduce visual artifacts at object boundaries. Figure S2, cropped from our ablation experiments (Section 5.1), highlights its significance in our method.
S3 Per-scene Results
Table S1, Table S2 and Table S3 present the per-scene values that are averaged in Table 3.
S4 Fully Solid Scenes
Although not our targeted use case, we tested our method on the fully solid scenes of the NeRF-Synthetic dataset [34], which lacks fuzzy objects. As noted in Section 5.2, our advantages in these scenes are marginal. While we outperform PermutoSDF [44], our quality remains behind other baselines. We model fuzzy surfaces by optimizing sample distribution rather than reconstructing high-frequency geometric details, as spatial sampling is key to accurately capturing these effects. By favoring smoother surfaces, our method tends to reconstruct overly simplified geometry in under-observed areas (Figure S7). Fully solid scenes can be optimally modeled as a single surface. However, SDF-based methods struggle in handling thin structures, as optimization often fails to reconstruct them reliably (e.g., BakedSDF [60], BOG [43]). Our surface smoothing, combined with view-dependent transparency, often leads to thin structures being reconstructed as view-dependent effects (Figure S8). As a result, our model tends to overfit training views, leading to a larger quality gap between training and test views (Table S4).
Refer to caption Refer to caption Refer to caption Refer to caption
PermutoSDF 7-Mesh (ours) PermutoSDF 7-Mesh (ours)
Refer to caption Refer to caption Refer to caption Refer to caption
3DGS Ground truth 3DGS Ground truth
Figure S4: Qualitative comparison of our 7-Mesh model with PermutoSDF [44] and 3DGS [24]. Scenes from the DTU dataset [21].
Refer to caption Refer to caption
$a$) 3DGS $b$) 7-Mesh (ours)
Refer to caption Refer to caption
$c$) 3DGS-75K $d$) Ground truth
Figure S5: 3DGS [24] demonstrates superior performance in modeling thin structures but is significantly less effective in representing large, textured areas. Our method renders faster than 3DGS-75K on mobile devices. Results on the khady scene from Shelly [56]. Quantitative results are in Table 2.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
$a$) 5-Mesh (ours) $b$) MobileNeRF $c$) Ground truth
Figure S6: Our method surpasses MobileNeRF [7] in modeling volumetric hair while also achieving superior performance on flat surfaces. Results on hairy monkey from QuadFields [47], our custom plushy scene and khady from Shelly [56]. Quantitative results are in Table 2.
Figure S7: Visualization of surface normals from our 7-Mesh model. Scene from NeRF-Synthetic [34].
Table S1: Per-scene results for baselines. Refer to Table 3 for averaged results. Metrics not provided are denoted with “—”. PermutoSDF trained until densities are fully peaked ($\phi_{\beta_{3}})$. The hairy monkey scene is from Sharma et al. [47].
PermutoSDF [44] 3DGS [24] MobileNeRF [7]
Dataset Scene PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Shelly [56] fernvase 28.42 0.953 0.078 34.82 0.986 0.040 29.06 0.957 0.088
horse 34.68 0.993 0.040 41.45 0.997 0.038 33.31 0.988 0.065
khady 26.22 0.879 0.226 30.54 0.924 0.187 26.42 0.877 0.228
kitten 30.91 0.971 0.093 38.17 0.991 0.050 30.22 0.968 0.098
pug 29.48 0.953 0.168 35.96 0.983 0.089 28.57 0.927 0.197
woolly 29.39 0.949 0.167 31.71 0.969 0.130 28.20 0.919 0.221
Custom hairy monkey 33.67 0.977 0.194 37.67 0.990 0.142 30.25 0.949 0.200
plushy 32.94 0.945 0.192 37.02 0.975 0.153 31.53 0.934 0.190
DTU [21] scan 105 34.78 0.985 0.124 35.50 0.984 0.102
scan 83 37.84 0.991 0.072 40.61 0.994 0.070
Table S2: Our per-scene results. The hairy monkey scene is from Sharma et al. [47]. Refer to Table 3 for averaged results.
@@
Shelly [56] fernvase 32.41 0.985 0.066 33.63 0.988 0.064
horse 38.34 0.998 0.038 39.78 0.998 0.034
khady 29.78 0.938 0.193 29.88 0.941 0.194
kitten 35.84 0.991 0.078 36.85 0.992 0.076
pug 33.72 0.983 0.138 34.25 0.985 0.132
woolly 30.26 0.973 0.175 31.12 0.978 0.162
Custom hairy monkey 35.59 0.984 0.178 35.90 0.985 0.179
plushy 34.41 0.957 0.164 34.99 0.965 0.164
DTU [21] scan 105 34.77 0.980 0.120 35.40 0.982 0.106
scan 83 38.05 0.990 0.064 38.34 0.990 0.063
Table S3: Our per-scene results. The hairy monkey scene is from Sharma et al. [47]. Refer to Table 3 for averaged results.
@@
Shelly [56] fernvase 34.55 0.990 0.062 34.64 0.991 0.062
horse 40.05 0.998 0.033 39.32 0.998 0.034
khady 29.97 0.942 0.194 29.96 0.943 0.195
kitten 37.11 0.993 0.074 37.05 0.993 0.074
pug 34.25 0.985 0.132 34.24 0.985 0.133
woolly 31.04 0.978 0.158 31.05 0.978 0.160
Custom hairy monkey 36.09 0.987 0.177 36.14 0.987 0.175
plushy 35.18 0.967 0.162 35.35 0.969 0.160
DTU [21] scan 105 35.50 0.982 0.106 35.54 0.983 0.105
scan 83 38.04 0.991 0.062 38.81 0.991 0.062
Table S4: Results averaged across test scenes. In solid scenes, our method outperforms PermutoSDF [44] (see Figure S8) but lags behind other baselines. We explain this behavior in Section S4. Methods marked with a $\star$ show results taken from original papers. PermutoSDF trained until densities are fully peaked ($\phi_{\beta_{3}})$. Metrics not provided by a baseline are denoted with “—”.
@@
3DGS [24] 36.76 0.991 0.030 33.23 0.981 0.037
Instant-NGP [35] 33.18
PermutoSDF [44] 29.31 0.975 0.057 28.05 0.966 0.065
AdaptiveShells [56] 31.84 0.957 0.056
QuadFields [47] 31.00 0.952 0.069
MobileNeRF [7] 30.90 0.947 0.060
3-Mesh 32.40 0.983 0.060 28.50 0.958 0.083
5-Mesh 33.23 0.986 0.055 28.77 0.959 0.081
7-Mesh 33.31 0.986 0.056 28.88 0.960 0.081
9-Mesh 33.19 0.986 0.057 28.79 0.960 0.082
Refer to caption Refer to caption Refer to caption Refer to caption
$a$) PermutoSDF $b$) 7-Mesh (ours) $c$) PermutoSDF $d$) 7-Mesh (ours)
Figure S8: Qualitative comparison of our method with PermutoSDF [44]. Renderings ($a$) and ($b$) are from a training view. Renderings ($c$) and ($d$) are from a test view. Scene from NeRF-Synthetic [34].